SUB-DOPPLER HIGH RESOLUTION EXCITATION SPECTROSCOPY OF DIBENZOFURAN

\underline{MASAAKI BABA}, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; AT-SUSHI DOI, YOSHIO TATAMITANI, MICHIRU YAMAWAKI, SHUNJI KASAHARA, and HA-JIME KATÔ, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan.

Dibenzofuran (DF) is one of the prototypical molecules of toxic dioxins as well as dibenzo-p-dioxin.^{a} Sub-Doppler highresolution fluorescence excitation spectra of the S_{1} ←S_{0} transition of DF have been observed in a collimated molecular beam. Each rotational line was fully resolved and the rotational constants in both the S_{0} and S_{1} states were determined by the rotational analysis of the 0-0 band (33773.478 cm^{−1}). The DF molecule has shown to be planar both in the S_{0} and S_{1} states. All of the observed lines were the ΔK = ±1 transition (B type). Therefore, the S_{1} state is assigned to be the ^{1}A_{1} state in C_{2v} symmetry.

aM. Baba, A. Doi, Y. Tatamitani, S. Kasahara, and H. Katô, J. Phys. Chem. A, 108, 1388 (2004)